A gold chain is sold for $ 635 at a gain of 27 % on the sale price. Find the profit.
Found 2 solutions by nerdybill, Alan3354:
Answer by nerdybill(7384)

A gold chain is sold for $ 635 at a gain of 27 % on the sale price. Find the profit.
.
.27 * 635 = $171.45

Answer by Alan3354(69443)

$635 = 1.27*W (W = wholesale)
W = 635/1.27 = $500
Profit = 635 - 500 = $135